# Benchmarking Multi-Modal Retrieval for Long Documents

### v1

@@ Abstract.
Multi-modal document retrieval is designed to identify and retrieve various forms of multi-modal content, such as figures, tables, charts, and layout information from extensive documents. Despite its significance, there is a notable lack of a robust benchmark to effectively evaluate the performance of systems in multi-modal document retrieval. To address this gap, this work introduces a new benchmark, named as MMDocIR, encompassing two distinct tasks: page-level and layout-level retrieval. The former focuses on localizing the most relevant pages within a long document, while the latter targets the detection of specific layouts, offering a more fine-grained granularity than whole-page analysis. A layout can refer to a variety of elements such as textual paragraphs, equations, figures, tables, or charts. The MMDocIR benchmark comprises a rich dataset featuring expertly annotated labels for 1,685 questions and bootstrapped labels for 173,843 questions, making it a pivotal resource for advancing multi-modal document retrieval for both training and evaluation. Through rigorous experiments, we reveal that (i) visual retrievers significantly outperform their text counterparts, (ii) MMDocIR train set can effectively benefit the training process of multi-modal document retrieval and (iii) text retrievers leveraging on VLM-text perform much better than those using OCR-text. These findings underscores the potential advantages of integrating visual elements for multi-modal document retrieval. MMDocIR is available at https://huggingface.co/MMDocIR.
Multi-modal Document Retrieval, Page Retrieval, Layout Retrieval
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: ; ; isbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Multi-modal Document Retrieval
@@
Figure 1. MMDocIR comprises 313 lengthy documents across 10 different domains, along with 1,685 questions. For each question, page-level annotations are provided via selected screenshots. Red boundary boxes represent layout-level annotations.
@@
Figure 2. Area ratio of different modalities (1) in overall and (2) by domains in MMLongBench-Doc benchmark (Ma et al., 2024b). Note that the white spaces, headers, and footers are removed from the area counting.
1. Introduction
Multi-modal document retrieval (Hassan et al., 2013; Ma et al., 2024a; Faysse et al., 2024; Lee et al., 2024) aims to encode and retrieve various forms of content from visually rich documents based on user queries. Unlike traditional document retrieval (Zhang et al., 2022; Chen et al., 2023; Sarthi et al., 2024; Dong et al., 2024) which primarily deals with textual data, multi-modal document retrieval must handle documents that seamlessly integrate text with multi-modal elements such as images, tables, charts, and layout designs. These multi-modal content often carry significant information that plain text cannot convey (Cui et al., 2021; Zhu et al., 2024; Sassioui et al., 2023), thereby enriching the depth and relevance of retrieved content. For example, our analysis of MMLongBench-Doc benchmark (Ma et al., 2024b) reveals that: text occupies only 52.7% of content area, while images and tables account for 29.2% and 12.8% respectively, as shown in Figure 2. This significant presence of multiple data modalities requires retriever to leverage all data modalities present in a document.
However, current benchmarks (shown in Table 1) for evaluating multi-modal document retrievers are insufficient, lacking in certain aspects that are critical for a comprehensive assessment. The major shortcomings include: 1. Question Quality: The design and curation of questions in most benchmarks do not align with the specific needs of multi-modal document retrieval. The questions, often reused from Visual Question Answering (VQA) tasks, lack the specificity required for this purpose and require expert filtering to be truly effective for retrieval tasks. 2. Document Quality: A significant limitation of existing benchmarks is the missing of complete document pages, thus hindering accurate evaluation of retrieval capabilities within actual document contexts. Additionally, the range of document domains covered is often too narrow, limiting the benchmarks’ applicability across different fields. 3. Retrieval Granularity: Most benchmarks, except for SciMMIR (Wu et al., 2024), only allow retrieval at a page level which is insufficient, as user queries may pertain to specific parts of a page like a particular figure or table rather than the entire page.
Recognizing these gaps, we introduce MMDocIR Multi-Modal Document Information Retrieval benchmark. MMDocIR is innovatively structured around two critical tasks: page-level and layout-level retrieval. (1) The page-level retrieval task is designed to identify the most relevant pages within a document in response to a user query. (2) The layout-level retrieval aims to retrieve most relevant layouts. The layouts are defined as the fine-grained elements such as paragraphs, equations, figures, tables, and charts. This task allows for a more nuanced content retrieval, honing in on specific information that directly answers user queries. To support these tasks, we first develop MMDocIR evaluation set comprising 313 documents averaging 65.1 pages each, along with VQA 1,658 questions derived from MMLongBench-Doc (Ma et al., 2024b) and DocBench (Zou et al., 2024). Initially, we examine all VQA questions, filter out questions that are irrelevant to IR tasks, and revise remaining questions to be more suitable for IR. Then, we manually annotate these questions with page and layout labels. The page-level labels are the specific pages identified that contain answer evidence.  [^1] It is worth noting that while MMLongBench-Doc provided initial page labels, our meticulous review lead to corrections in about 21.3% of these labels. The layout-level labels are the bounding boxes meticulously delineated around the key evidence within the identified pages. Meanwhile, we introduce MMDocIR training set, comprising 73,843 QA pairs converted from 7 DocQA datasets. Specifically, we manually collect 6,878 documents and use either manual or automatic methods to annotate the ground truth labels.  [^2] Since layout annotations are missing for most datasets, our data annotating pipeline can only create layout-level labels for 4 subsets.
@@ 1. Introduction
Type By Expert? For IR? #Num Evidence Type Domain #Pages Source Page Layout
DocCVQA (Tito et al., 2021) VQA question 20 TXT/L Finance 1.0
SciMMIR (Wu et al., 2024) Image caption 530K TAB/I Science 1.0
ViDoRe (Faysse et al., 2024) VQA question 3,810 TXT/C/TAB/I Multi-domain 1.0
PDF-MVQA (Ding et al., 2024) Search query 260k TXT/TAB/I Biomedical 9.6
MMLongBench-Doc (Ma et al., 2024b) VQA question 1,082 TXT/L/C/TAB/I Multi-domain 47.5
Wiki-SS (Ma et al., 2024a) Natural question 3,610 TXT Wikipedia 1.0
DocMatix-IR (Ma et al., 2024a) VQA question 5.61M TXT/L/C/TAB/I Multi-domain 4.2
MMDocIR VQA question 1,658 TXT/C/TAB/I Multi-domain 65.1
Table 1. Comparison between our benchmark and previous Document IR datasets. TXT/L/C/TAB/I: pure text/generalized layout/chart/table/image.
@@
In our research, we conduct a thorough evaluation of the prevailing multi-modal document retrieval baselines which can broadly be categorized into visual-driven and text-driven retrievers. Visual-driven retrievers (Ma et al., 2024a; Faysse et al., 2024), leverage vision-language models (VLMs) (Abdin et al., 2024; Beyer et al., 2024; Bai et al., 2023; Chen et al., 2024b) to capture rich multi-modal cues and generate embeddings for both queries and documents. In contrast, text-driven retrievers (Karpukhin et al., 2020; Khattab and Zaharia, 2020; Izacard et al., 2022; Wang et al., 2022; Xiao et al., 2023; Li et al., 2023) primarily rely on OCR or VLM to first convert the multi-modal content into text, subsequently employing language models (LMs)  (Devlin et al., 2019; Liu et al., 2019; Wang et al., 2024b) to generate embeddings for both queries and documents. Via extensive experiments, we discover that visual-driven retrievers consistently outperform their text-driven counterparts, often by a significant margin. This highlights the critical role that visual elements play in multi-modal document retrieval. In summary, our contributions are three-fold:
@@
  • Dual-task Retrieval Framework: We propose a dual-task retrieval framework (§$\S$ 2) that encompasses page-level and fine-grained layout-level multi-modal document retrieval.
@@
  • MMDocIR Benchmark: We introduce Multi-Modal Document Information Retrieval (MMDocIR) benchmark. The evaluation set (§$\S$ 3) consists of 313 documents with expert-annotated labels for 1,658 QA pairs. The training set (§$\S$ 4) consists of 6,878 documents and labels for 73,843 QA pairs.
  • Comprehensive Evaluation: We conduct an extensive evaluation of existing retrieval systems (§$\S$ 6). Our results reveal that visual-driven retrievers consistently outperform their text-driven counterparts, highlighting the importance of utilizing visual information in multi-modal document retrieval.
2. Dual-Task Retrieval Definition
@@
Let $\mathcal{D}$ be a document corpora consisting of a set of document pages: $\mathcal{P}=\{p_{1},p_{2},\dots,p_{n}\}$, and layouts: =$\mathcal{L}=\{l_{1},l_{2},\dots,l_{m}\}$ extracted via layout detection [^3] A document page usually comprises about 5 to 15 layouts, depending on its complexity.. The objective is to perform document page-level and layout-level retrieval. Specifically, the task involves retrieving a subset of these pages and layouts that are most relevant to a query $Q$, identified by the top $k$ entries where $k$ is significantly smaller than $n$ or $m$. The relevance of pages ($p$) and layouts ($l$) to $Q$ is quantified by similarity scores, $\textrm{Sim}(Q,p)$ and $\textrm{Sim}(Q,l)$ respectively. The retrieval system can be decomposed into (1) an offline indexing phase in which each page and layout from $\mathcal{P}$ and $\mathcal{L}$ is transformed into a vector representation, and (2) an online querying phase in which a query $Q$ is converted into a vector form, and the vectorized query is then compared against the indexed vectors using the similarity scores $\textrm{Sim}(Q,p)$ for pages and $\textrm{Sim}(Q,l)$ for layouts.
3. MMDocIR: Evaluation Set
{tabu}
l—c@ c@ c@ c@ c@ c@ c—c@   c@  c@   c@   —c@   c@  c@   c@
Dataset Domain Document Statistics   Questions (%)   Modality Distribution (%)
@@ 3. MMDocIR: Evaluation Set
Table 2. Detailed statistics for MMDocIR evaluation set. “#Lay/Page” is the averaging layouts per page, reflecting page’s layout complexity. “%Lay” refers to the area ratio of useful layouts (excluding white spaces, headers, and footers) over entire page.
3.1. Document Corpora Collection
To facilitate the development of MMDocIR, we leverage resources from existing Document Visual Question Answering (DocVQA) benchmarks (Mathew et al., 2021; Tito et al., 2023; Tanaka et al., 2023; Zhu et al., 2022; Li et al., 2024; Landeghem et al., 2023; Ma et al., 2024b; Zou et al., 2024). Although DocVQA benchmarks are not primarily designed for IR, they offer valuable document corpora and questions that can be adapted for our purposes. To select useful DocVQA datasets, we adhere to following criteria:
  • Diverse Domain/Modality: The document collections must (1) encompass diverse domains e.g., academia, finance, publishing, and research, and (2) embrace multiple modalities, such as text, figures, tables, charts, and layouts.
  • Long Document: We seek documents of sufficient length as longer texts pose more significant challenges. This criterion can evaluate models in handling complex and lengthy documents.
  • QA Diversity and Comprehensiveness: The questions included in the dataset should be diverse and challenging. For example, some questions should require reasoning across both text and visual tables/figures, while some are multi-hop questions that require synthesizing information from multiple pages.
After considering aforementioned criteria, we choose to leverage on the document corpora and questions from MMLongBench-Doc (Ma et al., 2024b) and DocBench (Zou et al., 2024). MMLongBench-Doc is a long-context, multi-modal benchmark comprising 1,091 questions constructed upon 135 documents averaging 47.5 pages. Its corpora include domains of research report, administration and industry, tutorial and workshop, academic paper, brochure, financial report, and guidebook. DocBench focuses on long document understanding, consisting of 1,102 questions constructed upon 229 documents averaging 77.5 pages. The corpora span across domains: academia, finance, government, laws, and news. Both datasets provide expert-annotated questions, requiring content evidence of various modalities (e.g., text, tables, charts and images, as well as locations layouts) to answer. Consequently, we obtain a total of 364 documents and 2,193 questions for our subsequent annotation.
3.2. Annotation Process
Question Filtering and Revision
To ensure that the questions in MMDocIR are optimally suited for document retrieval tasks. To this end, we identify four specific types of questions that do not align well with the objectives of IR. By filtering out or modifying these questions, we ensure the integrity and relevance of MMDocIR.
  • Summarization Questions: These questions, such as “What does this book mainly illustrate?” require a comprehensive understanding of large sections of a document or even the entire document. The broad scope makes it hard to pinpoint specific content effectively, which contradicts the precise nature of IR tasks.
  • Overwhelm Statistical Questions: Questions that demand extensive data computation or collation, such as “How many words are there in total in the paper?” also fall outside our scope. These require a level of statistical analysis that goes beyond retrieval.
  • Online Search Questions: For example, queries like “What is the Google Scholar citation count of the main author of this article?” rely on information from external online resources. Since our focus is on retrieving information that is internally available within the documents, these questions are excluded.
Page-level annotation process
To address the challenge of accurately locating relevant information within extensive documents, we annotate page-level labels that pinpoint the exact pages containing evidence needed to answer specific questions. Considering that the average document in our dataset contains 65.1 pages, pinpointing the correct pages is similar to finding a needle in a haystack and requires meticulous effort. Our annotation process is:
  • For MMLongBench-Doc: we review and correct the answers and corresponding page labels for 794 questions. Specifically, we make corrections to 10 answers and revise 169 page labels [^4] Common errors in page labeling: annotators starting page indexing at 1, missing labels for questions spanning multiple pages, and incorrect or absent page labels. to ensure they accurately reflect the evidence locations.
  • For DocBench: we annotate page labels for all 864 questions from scratch. This involves a detailed examination of each document to accurately assign page labels that match the evidential requirements of each question.
Through these efforts and following a rigorous annotation and cross-validation process, we successfully compile page-level labels for a total of 1,658 questions in MMDocIR.
Layout-level annotation process
To achieve more fine-grained retrieval granularity, we go beyond page-level annotations by including layout-level labels. This advanced annotation process captures the specific layout elements within pages that contain necessary evidence to answer each question. Unlike page-level annotation, labeling layout-level information is more complex and labor-intensive due to its detailed nature. Our annotations are via:
  • To streamline annotation process, we utilize MinerU (Wang et al., 2024a) to parse all documents and extract the layout information (e.g., type and bounding boxes) for every layout element on a page. These elements are categorized into five main types: text, image, table, title, and equation. The bounding boxes are delineated by the coordinates of the top-right and bottom-left corners.
  • Based on the layouts drawn by MinerU, we identify the layouts that contain necessary eivdence to answer the question. In instances where MinerU fails to detect certain evidentiary elements, we manually draws the bounding boxes to ensure the precision of layout labels. This manual intervention is necessary for approximately 7% of our layout-level labels.
Ultimately, this meticulous process leads to the annotation of 2,638 layout labels for the 1,658 questions in MMDocIR.
3.3. Quality Control
Consistency Page Labels Layout Labels
Prec. Recall F1 Prec. Recall F1
[…]
Average 95.0 95.4 95.2 87.0 87.2 87.1
[…]
{tabu}
l—c—ccc@ c@ c—c@   c@  c@   c@   —c@   c@
Dataset Domain Document Statistics   Modality Distribution (%)   Labels
@@ 3.3. Quality Control
MMDocIR Train assorted docs 6,878 73,843 32.6 6.32 42.6 49.3 34.3 10.8 4.9
@@
- MP-DocVQA (Tito et al., 2023) health/ind. docs 875 15,266 46.8 6.9 38.8 57.3 18.0 22.7 1.9
- SlideVQA (Tanaka et al., 2023) diverse slides 2,011 11,066 49.3 4.4 42.3 30.1 56.2 4.7 8.8
- TAT-DQA (Zhu et al., 2022) annual reports 163 15,814 147.3 9.2 42.2 66.4 4.4 26.5 2.7
- ArXivQA (Li et al., 2024) arXiv papers 1,579 12,314 18.4 7.9 50.0 70.4 22.3 2.8 1.0
- SciQAG (Wan et al., 2024) science papers 1,197 4,976 9.0 9.1 53.7 61.8 28.0 6.7 1.5
- DUDE (Landeghem et al., 2023) assorted docs 779 3,173 15.6 7.4 42.5 57.1 24.7 15.2 2.9
- CUAD (Hendrycks et al., 2021) legal contract 274 11,234 29.6 7.4 24.7 89.3 2.5 6.4 1.1
Table 4. Document statistics for Training Datasets collected.
To ensure annotation quality and reliability in MMDocIR, we have adopted a rigorous quality control process using a cross-validation method. This method involves splitting the document corpora into two main groups, labeled as Group A and Group B. Each group is responsible for annotating page-level and layout-level labels for approximately 1,000 QA pairs. Importantly, there is an overlap of 400 QA pairs between two groups, which serves as a validation checkpoint. The process for quality control works as follows:
  • Overlap Scoring: For 400 overlapping questions, each group’s annotations are scored against the other’s as a form of mutual validation. Specifically, Group A’s annotations are used as a ground truth to score Group B’s work, and conversely, Group B’s annotations serve as the ground truth for scoring Group A.
  • Cross-Evaluation: By cross-evaluating the outcomes from both groups, we are able to assess the consistency of the annotations and identify any discrepancies for improvement. Specifically, we achieved an F1 score of 95.2 for page-level annotations and 87.1 for layout-level annotations, as shown in Table 3.
  • Random Cross-Validation: Apart from the overlapping 400 questions, we randomly cross-validate around 50% of the remaining annotations. In the cases where we have different opinions, we discuss to achieve mutually-agreed annotations. This step ensures the high quality and reliability of MMDocIR.
3.4. Statistics and Analysis

### v2

# MMDocIR: Benchmarking Multimodal Retrieval for Long Documents

@@ Abstract
Multimodal document retrieval aims to identify and retrieve various forms of multimodal content, such as figures, tables, charts, and layout information from extensive documents. Despite its increasing popularity, there is a notable lack of a comprehensive and robust benchmark to effectively evaluate the performance of systems in such tasks. To address this gap, this work introduces a new benchmark, named MMDocIR, that encompasses two distinct tasks: page-level and layout-level retrieval. The former evaluates the performance of identifying the most relevant pages within a long document, while the later assesses the ability of detecting specific layouts, providing a more fine-grained measure than whole-page analysis. A layout refers to a variety of elements, including textual paragraphs, equations, figures, tables, or charts. The MMDocIR benchmark comprises a rich dataset featuring 1,685 questions annotated by experts and 173,843 questions with bootstrapped labels, making it a valuable resource in multimodal document retrieval for both training and evaluation. Through rigorous experiments, we demonstrate that (i) visual retrievers significantly outperform their text counterparts, (ii) MMDocIR training set effectively enhances the performance of multimodal document retrieval and (iii) text retrievers leveraging VLM-text significantly outperforms retrievers relying on OCR-text. Our dataset is available at https://mmdocrag.github.io/MMDocIR/.
MMDocIR: Benchmarking Multimodal Retrieval for Long Documents
[…]
@@
Figure 1: MMDocIR evaluation set comprises 313 long documents and 1,658 queries across 10 domains. For each query, page-level labels are provided via selected screenshots. Red boundary boxes represent layout-level labels.
$\dagger$\$$\dagger$\$footnotetext: These authors contributed equally
@@
Figure 2: Area ratio of different modalities (1) in overall and (2) by domains in MMLongBench-Doc benchmark.
1 Introduction
@@
Type Expert? IR? #Num Evidence Type Domain #Pages Source Page Layout
DocCVQA VQA question 20 TXT/L Finance 1.0
SciMMIR Image caption 530k TAB/I Science 1.0
ViDoRe VQA question 3,810 TXT/C/TAB/I Multiple 1.0
PDF-MVQA Search query 260k TXT/TAB/I Biomedical 9.6
MMLongBench-Doc VQA question 1,082 TXT/C/TAB/I Multiple 47.5
Wiki-SS Natural question 3,610 TXT Wikipedia 1.0
DocMatix-IR VQA question 5.61m TXT/C/TAB/I Multiple 4.2
MMDocIR (eval) VQA question 1,658 TXT/C/TAB/I Multiple 65.1
MMDocIR (train) VQA question 73.8k TXT/C/TAB/I Multiple 49.3
Table 1: MMDocIR versus existing document IR datasets. TXT/C/TAB/I refers to text/chart/table/image.
Multimodal document retrieval Hassan et al. (2013); Lee et al. (2024) aims to retrieve information from visually rich documents based on user queries. Unlike traditional document retrieval Zhang et al. (2022); Chen et al. (2023) which primarily deals with textual data, multimodal document retrieval imposes substantially greater demands on understanding multimodal elements such as images, tables, charts, and layout designs. Such elements often carry significant information that plain text fails to convey Cui et al. (2021); Sassioui et al. (2023): tables reveal structured data patterns, charts visualize trends or correlations, images offer contextual and semantic cues, etc. Combining these visual elements enriches the quality of retrieved content. Our analysis of MMLongBench-Doc benchmark Ma et al. (2024b) in Figure 2 shows that: text occupies only 52.7% of content area, while images and tables account for 29.2% and 12.8% respectively. This highlights the need for retrieval systems that effectively handle multimodal information.
However, as shown on (see Table 1), existing benchmarks exhibit several critical limitations that undermine comprehensive evaluation of multimodal retrieval systems. The key limitations include: 1. Question Quality: Many questions used in existing benchmarks are directly sourced from datasets for Visual Question Answering (VQA) tasks. Some questions often assume the input is already relevant, making it not suited for meaningful evaluation of retrieval capabilities. 2. Document Completeness and Diversity: Existing benchmarks often provide only partial documents, limiting the ability to evaluate within full document context. Additionally, the narrow range of document domains further restricts their applicability across diverse use-cases in real-world. 3. Retrieval Granularity: Most benchmarks support only page-level retrieval. Such granularity is often insufficient, as user queries frequently target specific elements, such as figures or tables, rather than entire pages.
To address these gaps, we introduce MMDocIR, a multimodal document information retrieval benchmark. MMDocIR is designed for two key tasks: page-level and layout-level retrieval. (1) The page-level retrieval identifies the most relevant pages within a document to answer user query. (2) The layout-level retrieval targets the most relevant fine-grained elements such as paragraphs, equations, figures, tables, and charts (see Appendix E.2 for more examples). Such task supports more precise and context-aware retrieval that pinpoint specific elements to address user queries. To support both tasks, we develop MMDocIR evaluation set that comprises 313 documents, each averaging 65.1 pages, along with 1,658 modified queries derived from MMLongBench-Doc and DocBench Zou et al. (2024). The queries are annotated with 2,107 page-level and 2,638 layout-level labels. The page labels are specific pages that contain the evidence needed to answer the query. [^1] While MMLongBench-Doc provided initial page labels, our meticulous review lead to corrections in 21.3% of them. The layout labels consist of precisely drawn bounding boxes around the key evidence within the identified pages. In addition, we introduce the MMDocIR training set, designed to support retriever training. It contains 73,843 questions sourced from 7 DocQA datasets. To construct this set, we manually collect 6,878 documents and apply a semi-automatic pipeline to annotate the ground truth labels.
@@
By leveraging MMDocIR, we conduct a comprehensive evaluation on multimodal document retrieval across two retriever types: visual-driven and text-driven. Visual-driven retrievers Ma et al. (2024a); Faysse et al. (2024), leverage vision language models (VLMs) to capture rich multimodal cues and generate embeddings for both queries and documents. In contrast, text-driven retrievers Karpukhin et al. (2020); Khattab and Zaharia (2020); Xiao et al. (2023) rely on OCR or VLM to first convert the multimodal content into text, subsequently employing language models (LMs) to generate embeddings for both queries and documents. Our extensive experiments reveal that visual-driven retrievers consistently outperform their text-driven counterparts, often by a significant margin. In summary, our contributions are threefold:
@@
  • Dual-task Retrieval Framework: We propose a dual-task retrieval framework (§$\S$ 2) that supports page-level and fine-grained layout-level multimodal document retrieval.
@@
  • MMDocIR Benchmark: We introduce Multimodal Document Information Retrieval benchmark. The evaluation set (§$\S$ 3) consists of 313 documents with expert-annotated labels for 1,658 questions. The training set (§$\S$ 4) consists of 6,878 documents and labels for 73,843 questions.
  • We conduct extensive experiments and comparisons of both text and visual retrievers (§$\S$ 5), demonstrating clear advantage of incorporating visual content in multimodal retrieval tasks.
2 Dual-Task Retrieval Definition
@@
Let $\mathcal{D}$ be a document corpora consisting of document pages: $\mathcal{P}=\{p_{1},p_{2},\dots,p_{n}\}$, and layouts: =$\mathcal{L}=\{l_{1},l_{2},\dots,l_{m}\}$ extracted via layout detection. The objective is to perform document retrieval at both page-level and layout-level. Specifically, given query $Q$, the task is to retrieve the top $k$ pages and layouts most relevant to $Q$, where $k<<n$ and $k<<m$. The relevance of pages ($p$) and layouts ($l$) to $Q$ is measured by similarity scores, $\textrm{Sim}(Q,p)$ and $\textrm{Sim}(Q,l)$ respectively. The retrieval system consists of two phases: (1) an offline indexing phase, where pages and layouts from $\mathcal{P}$ and $\mathcal{L}$ are encoded into vectors, and (2) an online querying phase, in which a query $Q$ is encoded into a vector, which is then compared against the offline-indexed vectors using similarity scores $\textrm{Sim}(Q,p)$ for pages and $\textrm{Sim}(Q,l)$ for layouts.
3 MMDocIR: Evaluation Set
3.1 Document Corpora Collection
After a comprehensive review of existing DocVQA datasets, we select MMLongBench-Doc Ma et al. (2024b) and DocBench Zou et al. (2024) to facilitate our benchmark construction (see Appendix B.2 for our selection criteria). MMLongBench-Doc is a long-context, multimodal benchmark comprising 1,091 questions across 135 documents with 47.5 pages on average. DocBench emphasizes long document understanding, consisting of 1,102 questions across 229 documents, each with an average length of 77.5 pages. Both datasets offer corpora from diverse domains with expert-annotated questions that require evidence from various modalities. Consequently, we curate a set of 364 documents and 2,193 questions for our subsequent annotation.
3.2 Annotation Process
Question Filtering and Revision.
To ensure that the questions in MMDocIR are optimally suited for document retrieval tasks, we identify four specific types of questions (see Appendix B.3) that do not align well with the objectives of IR. By filtering and refining these questions, we ensure the integrity and relevance of MMDocIR, resulting in 1,658 questions for subsequent annotation.
Page-level Annotation.
We annotate page labels that precisely identify the exact pages containing ground truth evidence. Given that documents in MMDocIR contain 65.1 pages on average, pinpointing relevant pages is highly non-trivial, akin to finding a needle in haystack, which demands careful inspection and document understanding.
  • For DocBench: we manually annotate page labels for all 864 questions from scratch, by carefully reviewing each document and locating the pages containing answer evidence.
  • For MMLongBench-Doc: we rigorously review and validate the answers and page labels of 794 questions. This effort results in corrections to 10 answers and 169 page labels [^2] Common errors in page labeling: annotators starting page indexing at 1 rather than 0, missing labels for questions spanning multiple pages, and incorrect or absent page labels..
Layout-level Annotation.
To enhance the granularity of our benchmark, we extend our annotations to include layout-level labels, identifying specific layout elements as evidence. Compared to page annotation, layout-level labeling is significantly more complex and labor-intensive.
  • Layout Detection. We begin by utilizing MinerU Wang et al. (2024) to automatically parse all documents and detect all layouts (e.g., layout type and bounding boxes).
  • Evidence Identification. We identify the layouts that contain necessary answer evidence. In case where MinerU fails to detect evidentiary element, we manually annotate the the bounding boxes, accounting for 7% of the total layout-level labels.
3.3 Quality Control
Consistency Page Labels Layout Labels
Prec. Recall F1 Prec. Recall F1
[…]
Average 95.0 95.4 95.2 87.0 87.2 87.1
Table 2: Annotation consistency between group A & B.
To ensure annotation quality and reliability in MMDocIR, we have adopted a rigorous 3-stage quality control process. We split questions into two parts. Each group is responsible for annotating approximately 1,000 questions, with an overlap of 400 questions serving the need for cross-validation.
[…]
  • Cross-Evaluation: We cross-evaluate and achieve F1 score of 95.2 and 87.1 for page and layout labels, as shown in Table 2. We then identify and fix the discrepancies.
  • Random Cross-Validation: We randomly cross-validate 50% of the remaining annotations. In the cases where we have different opinions, we discuss to achieve mutually-agreed annotations.
(a) #words/layout
(b) Distribution density: #words/layout
(c) Distribution density: #words/page
Figure 3: Distribution of OCR/VLM-text by length.
3.4 Multimodal content as OCR/VLM-text
To apply multimodal retrieval to text retrievers, we convert multimodal layouts (e.g., tables or figures) into text. Specifically, we extract text using OCR Smith (2007) (“OCR-text”) and generate detailed descriptions using VLMs OpenAI (2024); Qwen-Team (2024) (“VLM-text”). As a result, each image layout is represented in three formats: original image, OCR-text, and VLM-text.
For layouts, the average word length and distribution of OCR-text and VLM-text of MMDocIR are shown in Figure 3(a) and 3(b). Notably, the length of VLM-text is 1.5 and 3.8 times of OCR-text for table (with more structured numbers) and figure (mostly with visual elaboration), respectively.
For pages, we construct two variants by combining the natural text with either OCR-text and VLM-text for each page, resulting in OCR-page and VLM-page representations. The average word length is 477 and 505 for OCR-page and VLM-page respectively, with their distribution shown in Figure 3(c).
MMDocIR Document Statistics Questions (%) Modality Distribution %
@@ 3.4 Multimodal content as OCR/VLM-text
Table 3: Detailed statistics of MMDocIR evaluation set. “#Lay/Page” is averaging layouts per page, reflecting layout complexity. “%Lay” is the area ratio of useful layouts (excluding white spaces, headers, and footers).
MMDocIR Domain Document Statistics Evidence Modality (%) Labels
@@
Train Subsets assorted docs 6,878 73,843 32.6 6.32 42.6 49.3 34.3 10.8 4.9
@@
- MP-DocVQA health/ind. docs 875 15,266 46.8 6.9 38.8 57.3 18.0 22.7 1.9
- SlideVQA diverse slides 2,011 11,066 49.3 4.4 42.3 30.1 56.2 4.7 8.8
- TAT-DQA annual reports 163 15,814 147.3 9.2 42.2 66.4 4.4 26.5 2.7
- arXivQA arXiv papers 1,579 12,314 18.4 7.9 50.0 70.4 22.3 2.8 1.0
- SciQAG science papers 1,197 4,976 9.0 9.1 53.7 61.8 28.0 6.7 1.5
- DUDE assorted docs 779 3,173 15.6 7.4 42.5 57.1 24.7 15.2 2.9
- CUAD legal contracts 274 11,234 29.6 7.4 24.7 89.3 2.5 6.4 1.1
Table 4: MMDocIR training set statistics about our collected documents, questions, and constructed labels.
3.5 Statistics and Analysis